# The emergence of sparse attention: impact of data distribution and benefits of repetition

### v1

@@ 3.2 The emergence of in-context learning depends on data as in the theory
Longer sequences and larger vocabulary size delay emergence. We begin by verifying that the intuition described above applies and that the learning curve exhibits the sharp phase transitions characteristic of emergence. This is indeed the case for sufficiently long sequences and/or large vocabularies, as illustrated in Figure 4. Importantly, there is only one phase transition despite two attention layers being needed to learn sparse patterns; Figure 17 shows that they are learned simultaneously. This observation is consistent with the results of Reddy [36] and Singh et al. [15].
@@
Repetition speeds up emergence but comes with overfitting risks. We next investigate the benefits of in-context and cross-sample repetition. From the high sensitivity of the emergence time on sequence length revealed by previous analysis, we expect in-context repetition to have a stronger effect than the cross-sample one. This is what we observe. The results reported in Figure 5, which evaluate the model performance on test data that has no repetition, demonstrate that repetition can significantly speed up emergence, by a factor of $4$ for in-context repetition and a factor of $2$ for cross-sample repetition. The attention sparsity perspective thus explains why in-context repetition has been found to favor in-context learning vs. in-weight learning [13, 36, 45]: the induction head needed to solve this kind of task is formed earlier with such a form of repetition.
However, this benefit comes with an overfitting risk: while repetition consistently accelerates learning (repetition always speeds up learning on the train loss, cf. Figure 16 in the appendix), too much repetition leads to learning strategies that do not generalize well. For example, selecting the most frequent value is a valid strategy for this task whenever the query appears two or more times in the context. Interestingly, we also observe some grokking-like patterns [46] as the test accuracy eventually starts to increase late in training for large amounts of repetition. We argue that this occurs because the no-repetition case is still represented in the training data, albeit with very little weight. This trade-off between learning speed and generalization is consistent with what Park et al. [31] reported on another in-context learning task. The overfitting we observe here may appear inconsistent with our theory but is not: our theory addresses learning speed (performance on the training loss) rather than generalization ability.
@@ 4 Discussion
Through the lens of sparse attention, we provide a unifying perspective on a set of empirical results highlighted in the Motivation (Section 1). While our focus on sparse attention and the effect of repetition is unique (to the best of our knowledge), there are multiple closely related works. First, the interaction between the feedforward weights and attention we study is closely related to the one between two consecutive feedforward linear layers [29, 47] [29, 47, 48]. Linear networks also display incremental learning with abrupt transitions characterizing each phase change [49, 50, 51, 52]. This line of research on linear networks has since then been extended to linear Transformers [45, 53], in particular to study the emergence of in-context learning. On the more conceptual side, there exist other theories of how emergence could arise from longer training, such as grokking [46, 14] or singular learning theory [54, 55]. While rarer, other theories focus on emergence from increased model size [56]. Compared to these, our theory focuses on learning dynamics (emergence over the course of training) and is directly tied to the internal mechanisms of the Transformers.
@@
One key result of our paper is that the learning of sparse attention is prone to producing emergence over the course of training. Given that sparse or concentrated attention is a common emergent feature of Transformers (e.g., induction heads [12], task/function vectors [57, 58], or high-norm tokens in vision transformers [59]), one may ask how much currently known emergent behaviors can be attributed to the learning of sparse attention patterns, and whether there are many more emergent behaviors than what is currently reported. However, these points must be weighed. As noted above, even simpler MLPs can give rise to abrupt emergence over training [e.g. 29]; thus, learning in other circuits within Transformers might contribute to sudden transitions in their performance. Furthermore, emergence results at large scale mostly show a sharp phase transition of the model as the number of FLOPs [e.g., 8], which roughly corresponds to the model size times the number of training steps, reaches a certain threshold. It is therefore unclear whether these examples of emergence are rooted in longer training, models with higher capacity, or a complex interplay between the two. More empirical and theoretical work is needed to better understand the causes of emergence in large models and the possible complex relationship between training-based and size-based emergence.
@@
Another key result of our work is that low data diversity can actually improve performance. This contrasts with classic machine learning principles, which state that low diversity hurts generalization. These two apparently conflicting statements are actually compatible. As our results highlight, low diversity accelerates the learning of sparse attention, but high data diversity is better as training time goes to infinity. We hypothesize that data diversity might be a powerful lever towards enabling active learning [60], with the ultimate goal of letting the learner decide what it wants to learn on and reaching human-level sample efficiency that current deep learning systems crucially lack [61, 62, 63]. Humans also encounter varying levels of data diversity over development; e.g., infants receive progressively increasing data diversity during their early years [64]. Such principles could already be at play in the current training pipeline of large language models, for instance when code, which typically has lower syntactic diversity than other types of text [65], is included at specific moments during training. While promising, this thread requires more extensive analysis, both at a larger scale and on more realistic data distributions.
@@ A.1 Overview of the assumptions
    Gradient flow dynamics. To study learning dynamics, we focus on the gradient flow of the expected loss. Compared to the standard deep learning regime, this removes stochastic noise induced by sampling, uses gradient descent as optimizer instead of optimizer with adaptive learning rates such as Adam, and requires an infinitely small learning rate. Due to the last point, phenomena such as the edge of stability [66] are out of the picture. Yet, this is a very standard assumption (e.g. [29]) and will enable us to use tools from dynamical systems to understand how the behavior of the model changes over the course of learning.
@@
    Initialize the weights $W$ at 0. When drawing the entries of $W$ i.i.d. and independently of those of $W^{*}$, $W$ is almost surely orthogonal to $W^{*}$ as $d\rightarrow\infty$. The components of $W$ not aligned with $W^{*}$ exhibit a fast decay towards 00, so this assumption corresponds to having already completed this process. With such an assumption, we can reduce the dynamics of the weights to a single scalar, $w$, which will be the projection of $W$ on a normalized $W^{*}$.
@@ A.2 Derivation of the expected loss and its gradients
In the third line, we used the fact that the first $T-B$ tokens are independent of each other and independent of the last $B$ ones, in the fourth line that the last $B$ tokens are all equal to $x_{T}$ and in the fifth line the equality $\mathbb{E}_{x}[\lVert Ax\rVert]=\mathrm{tr}(A\mathbb{E}[xx^{\top}]A^{\top})$.
@@
[…]
@@ A.3.1 Reducing the learning dynamics to two dimensions
  • Attention is initialized uniformly. The first assumption we make is that attention is perfectly uniform (Assumption 2 from Section A.1), that is $\alpha_{t}=\frac{1}{T}$ or $a_{t}=0$. Given that the gradients of these parameters are the same for $t\leq T-B$ and $t>T-B$, cf. the calculation in the previous section, all attention values will have two possible values. If we additionally leverage the fact that $\sum_{t}\alpha_{t}=1$, everything is captured by a single scalar $\Delta a$, that is defined as $a_{T}-a_{1}$. Indeed
@@
  • Weights are initialized at 00. We assume that $W=0$ at initialization, following Assumption 4 of Section A.1. As $\Sigma_{T}=\frac{1}{d}\mathrm{Id}$, we have
@@ A.3.2 Approximate dynamics around initialization
The Hessian matrix has a positive eigenvalue $\frac{B}{\sqrt{d}T}$ and one negative one $-\frac{B}{\sqrt{d}T}$. This implies that the initial parameters are in the vicinity of an unstable fixed point and that the negative eigenvalue of the Hessian will dictate how fast we are escaping these initial conditions. Its eigenvalues are
@@ A.3.3 Phase transition and late phase dynamics
This loss features multiplicative interactions akin to a one-hidden-layer neural network, and therefore it comes as no surprise that the loss exhibits similar sharp phase transitions as the ones observed when learning these networks, cf. Saxe et al. [29] for an in depth analysis of these dynamics (one needs to linearize the mapping $\Delta a\mapsto\alpha$ to get the exact mapping to this set of results).
@@ A.3.4 Empirical validation
Figure 9: The gradient flow dynamics on the simplified loss of Equation 20 match the stochastic gradient descent dynamics of the original objective. We report the dynamics of $5$ different seeds for the simulation lines. Changing the seed modified the target mapping, the sampling process, as well as the model initialization. Additional details are provided in Section A.3.4.
@@ A.4.1 Reducing the learning dynamics to three dimensions
Importantly, when initialized at 00, each column of $W$ will evolve on the line spanned by the corresponding column of $W^{*}$. However, they will not move at the same speed as the first column has different dynamics than the other columns. We therefore introduce two scalars $\tilde{w}$ and $w$ such that $W_{:1}=\tilde{w}W^{*}_{:1}$ and $W_{:i}=\frac{w}{\sqrt{d-1}}W_{:i}^{*}$. As for the analysis in the previous section, the $\sqrt{d-1}$ scaling factor is here to ensure that the learning speed coincide. To keep the notation as concise as possible, we introduce $\tilde{\sigma}:=p+\frac{1-p}{d}$ and $\sigma:=\frac{1-p}{d}$.
@@
$\frac{\mathrm{d}^{2}L}{\mathrm{d}\tilde{w}\,\mathrm{d}\Delta a}$ $=\alpha(1-\alpha)\left(-\frac{(1-\alpha)\tilde{w}}{d(T-1)}+\left(2\alpha\tilde{w}-1\right)\tilde{\sigma}\right)$
@@ A.4.2 Approximate dynamics around initialization
which means that the eigenvalues of the Hessian are 00 and
±$\pm\frac{1}{\sqrt{d}T}\sqrt{p^{2}d+(1-p)^{2}},$
@@ C.1 Implementation of the task
  • Set up query distribution. We define a query distribution $p_{\mathrm{query}}$ which corresponds to uniformly sampling one of the $2$-repeated tokens with probability $p$ and uniformly sampling all the tokens with probability $1-p$ (recall that $p$ is the cross-sample repetition probability). Said otherwise, this probability distribution has mass $\frac{p}{2}+\frac{1-p}{N_{\mathrm{tokens}}}$ on the repeated tokens and mass $\frac{1-p}{N_{\mathrm{tokens}}}$ on the rest of the tokens.
@@ D.2 Fitting power laws on plateau length
  • Figure 1.c: $C$, $\alpha$ and $\beta$ ($B$ is fixed to $1$).
@@ D.3 Architectural and training details
It should be additionally noted that we adapt the number of iterations depending on the difficulty of the task and roughly aim to end training significantly after emergence. However, for the most challenging versions of the task (e.g., large $T$ values in Figure 3), the phase transition sometimes occurs after the maximum number of training steps we attribute to the task (in that case $50$k steps).
@@
Layers $1$ $2$ $4$
Number of heads N/A $4$ $4$
@@ D.4 Reproducibility
Our theoretical simulations were run in JAX [67] and our Transformer experiments in PyTorch [68]. All our experiments were run on Nvidia RTX 3090 GPUs. The typical training run takes one hour, although training on shorter sequences can be significantly shorter. The code is publicly available at the following url: https://github.com/NicolasZucchet/The-emergence-of-sparse-attention.

### v2

@@ 3.2 The emergence of in-context learning depends on data as in the theory
Longer sequences and larger vocabulary size delay emergence. We begin by verifying that the intuition described above applies and that the learning curve exhibits the sharp phase transitions characteristic of emergence. This is indeed the case for sufficiently long sequences and/or large vocabularies, as illustrated in Figure 4. Importantly, there is only one phase transition despite two attention layers being needed to learn sparse patterns; Figure 17 shows that they are learned simultaneously. This observation is consistent with the results of reddy_mechanistic_2024 and singh_what_2024.
@@
Repetition speeds up emergence but comes with overfitting risks. We next investigate the benefits of in-context and cross-sample repetition. From the high sensitivity of the emergence time on sequence length revealed by previous analysis, we expect in-context repetition to have a stronger effect than the cross-sample one. This is what we observe. The results reported in Figure 5, which evaluate the model performance on test data that has no repetition, demonstrate that repetition can significantly speed up emergence, by a factor of $4$ for in-context repetition and a factor of $2$ for cross-sample repetition. The attention sparsity perspective thus explains why in-context repetition has been found to favor in-context learning vs. in-weight learning [chan_data_2022, reddy_mechanistic_2024, singh_transient_2024]: the induction head needed to solve this kind of task is formed earlier with such a form of repetition.
However, this benefit comes with an overfitting risk: while repetition consistently accelerates learning (repetition always speeds up learning on the train loss, cf. Figure 16 in the appendix), too much repetition leads to learning strategies that do not generalize well. For example, selecting the most frequent value is a valid strategy for this task whenever the query appears two or more times in the context. Interestingly, we also observe some grokking-like patterns [power_grokking_2022] as the test accuracy eventually starts to increase late in training for large amounts of repetition. We argue that this occurs because the no-repetition case is still represented in the training data, albeit with very little weight. This trade-off between learning speed and generalization is consistent with what park_competition_2024 reported on another in-context learning task. The overfitting we observe here may appear inconsistent with our theory but is not: our theory addresses learning speed (performance on the training loss) rather than generalization ability.
@@ 4 Discussion
Through the lens of sparse attention, we provide a unifying perspective on a set of empirical results highlighted in the motivation (Section 1). While our focus on sparse attention and the effect of repetition is unique (to the best of our knowledge), there are multiple closely related works. First, the interaction between the feedforward weights and attention we study is closely related to the one between two consecutive feedforward linear layers [saxe_exact_2014, saxe_mathematical_2019, varre_spectral_2023]. Linear networks also display incremental learning with abrupt transitions characterizing each phase change [jacot_saddle--saddle_2021, abbe_merged-staircase_2022, pesme_saddle--saddle_2023, abbe_sgd_2023]. This line of research on linear networks has since then been extended to linear [singh_transient_2024, zhang_training_2025] and softmax [chen_training_2024] Transformers, in particular to study the emergence of in-context learning. On the more conceptual side, there exist other theories of how emergence could arise from longer training, such as grokking [power_grokking_2022, nanda_progress_2023] or singular learning theory [watanabe_algebraic_2009, hoogland_developmental_2024]. While rarer, other theories focus on emergence from increased model size [arora_theory_2023]. Compared to these, our theory focuses on learning dynamics (emergence over the course of training) and is directly tied to the internal mechanisms of the Transformers.
@@
We find that the learning of sparse attention is prone to producing emergence over the course of training. Given that sparse or concentrated attention is a common emergent feature of Transformers (e.g., induction heads [olsson_-context_2022], task/function vectors [hendel_-context_2023, todd_function_2024], or high-norm tokens in vision transformers [darcet_vision_2024]), one may ask how much currently known emergent behaviors can be attributed to the learning of sparse attention patterns, and whether there are many more emergent behaviors than what is currently reported. However, these points must be weighed. As noted above, even simpler MLPs can give rise to abrupt emergence over training [e.g. saxe_exact_2014]; thus, learning in other circuits within Transformers might contribute to sudden transitions in their performance. Furthermore, emergence results at large scale mostly show a sharp phase transition of the model as the number of FLOPs [e.g., wei_emergent_2022], which roughly corresponds to the model size times the number of training steps, reaches a certain threshold. It is therefore unclear whether these examples of emergence are rooted in longer training, models with higher capacity, or a complex interplay between the two. More empirical and theoretical work is needed to better understand the causes of emergence in large models and the possible complex relationship between training-based and size-based emergence.
When does repetition help?
Overall, we expect repetition to be beneficial when training on tasks prone to performance plateaus, as there is a clear benefit in trading off learning speed for generalization in these cases. Cross-sample repetition should be particularly helpful in tasks where learning feedforward transformations is challenging (high $d$ in our analysis). This includes learning many factual associations as in zucchet_how_2025 and the reasoning-heavy arithmetic tasks of charton_emergent_2024, which require learning complex non-linear transformations. In-context repetition should be most beneficial when learning from very long sequences, as it effectively reduces the sequence length the model must process. Repetition, both in-context and cross-sample, is a natural feature of language that may accelerate some parts of language learning. The principles we touch upon could already be more directly at play in the current training pipeline of large language models, for instance when code, which typically has lower syntactic diversity than other types of text [hindle_naturalness_2016], is included at specific moments during training.
@@
A key result of our work is that low data diversity can actually improve performance. This contrasts with classic machine learning principles, which state that low diversity hurts generalization. These two apparently conflicting statements are actually compatible. As our results highlight, low diversity initially accelerates the learning of sparse attention, but high data diversity is better as training time goes to infinity. We hypothesize that data diversity might be a powerful lever towards enabling active learning [settles_active_2009]. Specifically, our findings suggest a simple active learning algorithm: whenever an agent detects it is stuck on a task, it can decrease data diversity to accelerate learning, then increase diversity after the critical transition occurs to improve generalization. This dynamic adjustment of diversity provides a practical mechanism for agents to actively control their own learning trajectories, with the ultimate goal of letting the learner decide what it wants to learn on and reaching human-level sample efficiency that current deep learning systems crucially lack [lake_human-level_2015, mnih_human-level_2015, warstadt_findings_2025]. Humans also encounter varying levels of data diversity over development. For example, infants receive progressively increasing data diversity during their early years [smith_developing_2018]. Our theory suggests that it may be an important factor in accelerating our development. While promising, this thread requires more extensive analysis, both at larger scale and on more realistic data distributions.
@@ A.1 Overview of the assumptions
    Gradient flow dynamics. To study learning dynamics, we focus on the gradient flow of the expected loss. Compared to the standard deep learning regime, this removes stochastic noise induced by sampling, uses gradient descent as optimizer instead of optimizer with adaptive learning rates such as Adam, and requires an infinitely small learning rate. Due to the last point, phenomena such as the edge of stability [cohen_gradient_2021] are out of the picture. Yet, this is a very standard assumption (e.g. [saxe_exact_2014]) and will enable us to use tools from dynamical systems to understand how the behavior of the model changes over the course of learning.
@@
    Initialize the weights $W$ at 0. When drawing the entries of $W$ i.i.d. and independently of those of $^{*}, WW$ is almost surely orthogonal to $W^{*}$ as $d\rightarrow\infty$. The components of $W$ not aligned with $W^{*}$ exhibit a fast decay towards 0, so this assumption corresponds to having already completed this process. With such an assumption, we can reduce the dynamics of the weights to a single scalar, $w$, which will be the projection of $W$ on a normalized $W^{*}$.
@@ A.2 Derivation of the expected loss and its gradients
In the third line, we used the fact that the first TBT-B tokens are independent of each other and independent of the last $B$ ones, in the fourth line that the last $B$ tokens are all equal to $x_{T}$ and in the fifth line the equality $\mathbb{E}_{x}[\lVert Ax\rVert]=\mathrm{tr}(A\mathbb{E}[xx^{\top}]A^{\top})$.
@@
$\nabla_{\alpha_{t}}L$ $=\left\{\begin{aligned} &\mathrm{tr}\left(W\Sigma_{T}\left(\sum_{t>T-B}\alpha_{t}W-W^{*}\right)^{\top}\right)~~\mathrm{if}~t>T-B\\ &\frac{\alpha_{t}}{d}\lVert W\rVert^{2}~~\mathrm{otherwise}.\end{aligned}\right$. (16)
@@ A.3.1 Reducing the learning dynamics to two dimensions
  • Attention is initialized uniformly. The first assumption we make is that attention is perfectly uniform (Assumption 2 from Section A.1), that is $\alpha_{t}=\frac{1}{T}$ or $a_{t}=0$. Given that the gradients of these parameters are the same for $t\leq T-B$ and t>TBt>T-B, cf. the calculation in the previous section, all attention values will have two possible values. If we additionally leverage the fact that $\sum_{t}\alpha_{t}=1$, everything is captured by a single scalar $\Delta a$, that is defined as $a_{T}-a_{1}$. Indeed
@@
  • Weights are initialized at 0. We assume that $W=0$ at initialization, following Assumption 4 of Section A.1. As $\Sigma_{T}=\frac{1}{d}\mathrm{Id}$, we have
@@ A.3.2 Approximate dynamics around initialization
The Hessian matrix has a positive eigenvalue $\frac{B}{\sqrt{d}T}$ and one negative one $\frac{B}{\sqrt{d}T}$. This implies that the initial parameters are in the vicinity of an unstable fixed point and that the negative eigenvalue of the Hessian will dictate how fast we are escaping these initial conditions. Its eigenvalues are
@@ A.3.3 Phase transition and late phase dynamics
This loss features multiplicative interactions akin to a one-hidden-layer neural network, and therefore it comes as no surprise that the loss exhibits similar sharp phase transitions as the ones observed when learning these networks, cf. saxe_exact_2014 for an in depth analysis of these dynamics (one needs to linearize the mapping $\Delta a\mapsto\alpha$ to get the exact mapping to this set of results).
@@ A.3.4 Empirical validation
Figure 9: The gradient flow dynamics on the simplified loss of Equation 20 match the stochastic gradient descent dynamics of the original objective. We report the dynamics of 55 different seeds for the simulation lines. Changing the seed modified the target mapping, the sampling process, as well as the model initialization. Additional details are provided in Section A.3.4.
@@ A.4.1 Reducing the learning dynamics to three dimensions
Importantly, when initialized at 0, each column of $W$ will evolve on the line spanned by the corresponding column of $W^{*}$. However, they will not move at the same speed as the first column has different dynamics than the other columns. We therefore introduce two scalars $\tilde{w}$ and $w$ such that $W_{:1}=\tilde{w}W^{*}_{:1}$ and $W_{:i}=\frac{w}{\sqrt{d-1}}W_{:i}^{*}$. As for the analysis in the previous section, the $\sqrt{d-1}$ scaling factor is here to ensure that the learning speed coincide. To keep the notation as concise as possible, we introduce $\tilde{\sigma}:=p+\frac{1-p}{d}$ and $\sigma:=\frac{1-p}{d}$.
@@
[…]
@@ A.4.2 Approximate dynamics around initialization
which means that the eigenvalues of the Hessian are 0 and
$\pm\frac{1}{\sqrt{d}T}\sqrt{p^{2}d+(1-p)^{2}},$
@@ C.1 Implementation of the task
  • Set up query distribution. We define a query distribution $p_{\mathrm{query}}$ which corresponds to uniformly sampling one of the $2$-repeated tokens with probability $p$ and uniformly sampling all the tokens with probability 1p1-p (recall that $p$ is the cross-sample repetition probability). Said otherwise, this probability distribution has mass $\frac{p}{2}+\frac{1-p}{N_{\mathrm{tokens}}}$ on the repeated tokens and mass $\frac{1-p}{N_{\mathrm{tokens}}}$ on the rest of the tokens.
@@ D.2 Fitting power laws on plateau length
  • Figure 1.c: $C$, $\alpha$ and $\beta$ ($B$ is fixed to 11).
@@ D.3 Architectural and training details
It should be additionally noted that we adapt the number of iterations depending on the difficulty of the task and roughly aim to end training significantly after emergence. However, for the most challenging versions of the task (e.g., large $T$ values in Figure 3), the phase transition sometimes occurs after the maximum number of training steps we attribute to the task (in that case 5050k steps).
@@
Layers 11 22 44
Number of heads N/A 44 44
@@ D.4 Reproducibility
Our theoretical simulations were run in JAX [bradbury_jax_2018] and our Transformer experiments in PyTorch [paszke_pytorch_2019]. All our experiments were run on Nvidia RTX 3090 GPUs. The typical training run takes one hour, although training on shorter sequences can be significantly shorter. The code is publicly available at the following url: https://github.com/NicolasZucchet/The-emergence-of-sparse-attention.